# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ Appendix F Comparison with principal component analysis
Our approach to analysis sets itself apart from that of Feng & Tu [32] principally in the selection of the basis $\{\vec{p_{i}},,i=1,\dots,d\}$ used for examining the weights. While they employ the principal components of the weight series - the eigenvectors of $\vec{\Sigma}$ we use the eigenvectors of the hessian matrix $\vec{H}(\vec{\theta_{K}})$ computed at the beginning of the analysis period.
@@
The principal component basis, as used by Feng & Tu, holds a distinct advantage. For our analysis, we needed to eliminate the near-linear trajectory of the weights by deducting the mean velocity. However, in Feng & Tu’s analysis, this movement is automatically subsumed in the first principal component due to its pronounced variance. Hence, there’s no necessity for additional subtraction of this drift in the weight covariance eigenbasis.
@@
The artifact issue becomes evident in Figure 3, which displays synthetic data generated through stochastic gradient descent within an isotropic quadratic potential coupled with isotropic noise. With 2,500 dimensions, the model mirrors the scale of a layer in the fully connected neural network that Feng & Tu investigated. The weight series comprises 12,000 steps, which correspond to ten epochs of training this network. Analyzing this data with the weight covariance eigenbasis seemingly suggests anisotropic variance and correlation time. However, if the data is inspected without any basis change, both the variance and correlation time appear isotropically distributed as anticipated.
To navigate around this key issue associated with the eigenbasis of $\vec{\Sigma}$, we adopted the eigenvectors of the Hessian matrix. Unlike $\vec{\Sigma}$, the Hessian is not computed as an average over update steps but can, in theory, be precisely calculated for any given weight vector. Consequently, the Hessian matrix does not suffer from finite size effects. The difference between these two bases for actual data is visible in Figure 4. Here, we analyzed only the weights of the first convolutional layer of the LeNet from the main text to ensure comparability with Feng & Tu’s results. In this specific comparison, the network was trained without weight decay. Due to this and the fact that we are only investigating the weights of one layer, $\lambda_{\textrm{cross}}$ is significantly larger than all Hessian eigenvalues. As a result, when analyzing in the eigenbasis of the Hessian matrix related to this layer, both the variance and the correlation time align well with the prediction for smaller Hessian eigenvalues.
However, analyzing in the eigenbasis of the weight covariance matrix, the correlation time appears heavily dependent on the second derivative of the loss in the given direction. Additionally, the relationship between the weight variance and the second derivative shifts and more closely aligns with Feng & Tu’s results as the power law exponent is significantly larger than one. The first principal component, which Feng & Tu referred to as the drift mode, stands out due to its unusually long correlation time. This is to be expected, as this is the direction in which the weights are moving at an approximately constant velocity.
@@
Figure 4: Comparison of weight and velocity variances for all 450 weights of the first convolutional layer of the LeNet, as discussed in the main text, analyzed in two different bases. In order to facilitate a more directly comparable analysis to Feng & Tu [32], the network was trained without weight decay for this specific analysis and the analysis period was limited to 10 epochs, as opposed to the usual 20 epochs. The columns represent different bases: for the left column $\vec{p_{i}}$ are the eigenvectors of $\vec{\Sigma}$ and for the right column $\vec{p_{i}}$ are the eigenvectors of $\vec{H}$. The mean velocity was subtracted in the right column. The rows illustrate the weight and velocity variance, $\sigma_{\theta,i}^{2}=\vec{p_{i}}^{\top}\vec{\Sigma}\vec{p_{i}}$, $\sigma_{v,i}^{2}=\vec{p_{i}}^{\top}\vec{\Sigma_{v}}\vec{p_{i}}$ (top row), and the correlation time $\tau_{i}=2\sigma_{\theta,i}^{2}/\sigma_{v,i}^{2}$ (bottom row). The second derivative of the corresponding direction is depicted on the x-axis, $\lambda_{i}=\vec{p_{i}}^{\top}\vec{H}\vec{p_{i}}$. The top row solid lines indicate fit regions for a linear fit. For the $\vec{H}$ eigenbasis, the respective exponent of the power law relation is $1.018\pm 0.008$ for weight variance and $1.017\pm 0.002$ for velocity variance with a $2\sigma$-error. For the $\vec{\Sigma}$ eigenbasis, the corresponding exponent is $1.537\pm 0.012$ for weight variance and $1.134\pm 0.002$ for velocity variance.
Appendix G Hessian Eigenvalue Density
Figure 5: The distribution of the approximated 5,000 Hessian eigenvalues of the LeNet discussed in the main text. The inset shows that the smallest approximated eigenvalue has a magnitude of about 0.005.
Appendix H Drawing with replacement
@@
Figure 6 offers clear visual proof that when examples are selected with replacement, the previously noted anti-correlations within the SGD noise vanish. This observation confirms our hypothesis that the anti-correlations mentioned in the main text are indeed an outcome of the epoch-based learning technique. Consequently, we can predict that this change will influence the behaviour of the weight and velocity variance. As previously discussed, the theoretical results we have achieved for Hessian eigenvectors with eigenvalues exceeding $\lambda_{\textrm{cross}}$ conform to what one would predict in the absence of any correlation within the noise. Therefore, when examples are drawn with replacement, we anticipate the weight variance to be isotropic in all directions, while the velocity variance should remain unchanged.
@@
Upon reviewing Figure 7, it is clear that the velocity variance stays unchanged as predicted. However, while the weight variance remains constant for a broader subset of Hessian eigenvalues, it reduces for extremely small eigenvalues. Likewise, the correlation time is still limited for these minuscule Hessian eigenvalues. These deviations can be attributed to the finite time frame of the analysis period, comprising 20,000 update steps. This limited time window sets a cap on the maximum correlation time, consequently leading to a decreased weight variance for these small Hessian eigenvalues. Despite this, it is noteworthy that this maximum correlation time is still roughly one order of magnitude longer than the maximum correlation time induced by the correlations arising from the epoch-based learning approach.
@@
Appendix I Loss and Accuracy during Training
Figure 8: The evolution of the loss (left) and accuracy (right) during training of LeNet described in the main text. The statistics are shown for both training and test set. For the first 100 epochs, the exponential learning rate decay was used, and for the last 20 epochs, the learning rate was fixed at the final value of the exponential decay.
Appendix J Testing different Hyperparameters
In this section we examine the dependence of the theoretical predictions on the three hyperparameters learning rate $\eta$, momentum $\beta$ and batch size $S$. For this, we train the LeNet again for 100 epochs, using an exponential learning rate schedule that reduces the learning rate by a factor of $0.98$ every epoch and afterwards we perform the numerical analysis as described in the main text.
However, we now train the network several times, always varying one of the hyperparameters while keeping the other two fixed. If not varied, the momentum was set to 0.90 and the batch size was set to 64. To ensure that training is always successful and 100% training accuracy is achieved, the initial learning rate was set to 0.005 when the batch size is varied and to 0.02 when the momentum is varied. Five different values are examined for each hyperparameter. To investigate the dependencies on the learning rate, the values 0.005, 0.01, 0.02, 0.03, and 0.04 were used for training. For momentum, the values 0.00, 0.50, 0.75, 0.90, and 0.95 were examined, and for batch size, the values 32, 50, 64, 100, and 128 were examined.
@@ Appendix I Loss and Accuracy during Training
Figure 9 shows as an example the weight and velocity variances as well as the correlation times for different values of the batch size for one training seed each. It can be seen that the theory is not only valid for the hyperparameter combination from the previous section, but is also generally applicable for different hyperparameters. In particular, we see that there is still good agreement with the theory even if the strictly necessary condition for the theoretical derivation of the noise autocorrelation, that the number of batches per epoch $M=N/S$ is an integer, is not met.
@@
For the evaluation of the dependence of these variables on the hyperparameters, they were determined as follows for the various hyperparameter combinations using the data from the respective correlation time plot. For the maximum correlation time $\tau_{\textrm{SGD}}$, the average of all correlation times was taken for which the corresponding Hessian eigenvalue is smaller than the theoretical crossover value. However, the result for the numerically determined maximum correlation time is not significantly different when simply taking the average of all determined correlation times for a hyperparameter combination, since only very few Hessian eigenvalues are larger than the crossover value.
For the numerical determination of the crossover value $\lambda_{\rm cross}$, a linear function was first fitted to the correlation times of the 20 largest Hessian eigenvalues in the log-log plot of the correlation times against the Hessian eigenvalues. In this region of the first 20 values, the correlation time is always clearly dependent on the Hessian eigenvalue and does not yet belong to the region of constant correlation times. The intersection of the fitted line with the numerically determined maximum correlation time $\tau_{\textrm{SGD}}$ is then taken as the crossover value $\lambda_{\rm cross}$. If the numerically determined correlation times follow the theory exactly, then the correlation times determined in this way for $\tau_{\textrm{SGD}}$ and $\lambda_{\rm cross}$ would also follow the theory accurately.
And indeed, Figure 10 shows a good agreement between the theory and the numerically determined values, although it should be noted that the deviations are larger than the random fluctuations between the different seeds.
@@
Appendix K Test Accuracy and the Hessian Eigendecomposition of the Weight Vector
Recent research [38, 39] suggests that the characteristics of a minimum in a neural network’s loss landscape are largely determined by the large Hessian eigenvalues and their corresponding eigendirections. However, our findings indicate that flat directions, which correspond to smaller eigenvalues, should not be overlooked. When we decompose a network’s weight vector in the Hessian eigenbasis, it becomes evident that both the projections onto high curvature directions and those onto relatively flat directions are important.
In Figure 11, we demonstrate this with the LeNet model. We analyzed the model’s test accuracy in relation to the exclusion of projections onto Hessian eigenvectors, starting with those possessing the largest eigenvalues. After calculating the top 10,000 Hessian eigenvalues and eigenvectors out of a total of 137,000, we observed that omitting the projection onto the top 35 eigenvectors (those above the crossover value $\lambda_{\rm cross}$) resulted in a decline in test accuracy from 63% to 54%. Discarding the top 5,000 eigenvectors, as analyzed in our main study, further reduced accuracy to 44%. It was only after excluding around 7,000 eigenvectors that the accuracy plummeted to 10%, akin to random guessing. This underscores the significance of the weight vector’s orientation relative to flat Hessian eigendirections for maintaining test accuracy, highlighting the potential impact of weight fluctuations in these directions.
The pronounced accuracy drop observed after removing around 6,000 eigenvectors may be attributed to the unexpectedly large projection of the weight vector onto these eigenvectors, which is apparent from Fig. Figure 11.
Figure 11: The top panel shows the test accuracy of the LeNet discussed in the main text and how it changes as we cumulatively discard projections onto more and more Hessian eigenvectors, starting with those with the largest eigenvalues. There is a sharp drop in accuracy for the first few eigenvectors, and then only after more than 6,000 discarded projections out of a total of 137,000 does the accuracy drop rapidly down to 10%, which is equivalent to random guessing. For comparison, the middle panel shows the corresponding Hessian eigenvalues and compares them to the crossover value $\lambda_{\rm cross}$, where the lower than expected weight variances hold only for eigendirections with eigenvalues smaller than this crossover value. It is in the range of eigenvalues significantly smaller than the crossover value that the accuracy drops sharply to 10%. The bottom panel shows the decomposition of the weight vector $\bf\theta$ in the Hessian eigenbasis. One explanation for the rapid drop in accuracy could be that the projection of the weights onto the corresponding eigenvectors is considerably larger than one would expect if the weight vector was a random vector. The expected $2\sigma$ interval for a random vector is displayed in red.
Appendix L Different Network Architecture
To further confirm our theoretical predictions within trained networks, in this section we turn to a more modern architecture. Instead of the previously used LeNet network, we examine the ResNet-20 network [5]. It is a convolutional network with significantly more convolutional layers than LeNet. It also uses residual blocks with residual connections, which allows for deeper network structures. As our loss function, we again employed Cross Entropy, along with an L2 regularization with a prefactor of $10^{-4}$ and we did not use batch normalization. The number of layers, which is already indicated in the name with 20, is significantly higher than in the LeNet with just five layers. With approximately 272,000 parameters, the ResNet-20 also has significantly more parameters and the computational cost is significantly higher.
Therefore, in this section we limit ourselves to examining the weight and velocity variances in the subspace of the 400 largest Hessian eigenvalues and associated eigenvectors. The network was trained with SGD for 100 epochs using the same exponential learning rate schedule as before, with a learning rate of $5\cdot 10^{-3}$, a momentum parameter of $0.9$, and a minibatch size of $50$. This setup achieves 100% training accuracy and 73% testing accuracy. In Figure 12 one can observe a good agreement between the theory predictions for the variances and correlation times and the numerical observations.
Figure 12: Relationship between Hessian eigenvalues and the variances and correlation times for a ResNet-20 trained on CIFAR10. The mean velocity of the weight trajectory was subtracted. The solid lines in the left panel indicate the regions utilized for a linear fit. The exponents resulting from the power law relationship are $0.820\pm 0.099$ for weight variance and $0.965\pm 0.005$ for velocity variance, with a $2\sigma$-error. The analysis was performed for the 400 largest Hessian eigenvalues and corresponding eigendirections.
Appendix M Effects of Non-Commutativity

### v3

@@ G.2 Comparison with principal component analysis
Our approach to analysis sets itself apart from that of Feng & Tu [32] principally in the selection of the basis $\{\vec{p_{i}},\,i=1,\dots,d\}$ used for examining the weights. While they employ the principal components of the weight series - the eigenvectors of $\vec{\Sigma}$ - we use the eigenvectors of the Hessian matrix $\vec{H}(\vec{\theta_{K}})$ computed at the beginning of the analysis period.
@@
The principal component basis, as used by Feng & Tu, holds a distinct advantage in their unregularized setup without weight decay. The advantage is connected to a near-linear drift of the weights they observed, even toward the end of training, in a direction roughly proportional to the overall weight vector. This behavior is a known instability in networks employing ReLU activations and softmax output layers: once 100% training accuracy is achieved, scaling all parameters simply scales the logits, thereby increasing the confidence of correct predictions without affecting classification outcomes. Applying sufficient regularization through weight decay eliminates this drift.
To ensure comparability with Feng & Tu, we omit weight decay in this appendix section. As a result, the weights exhibit a near-linear drift over time, similar to their observations. To isolate the stationary components of the dynamics, we subtract the mean velocity from both the weights and updates, redefining $\vec{\theta}_{k}$ and $\vec{v}_{k}$ as $\vec{\theta}_{k}^{\textrm{(s)}}\coloneqq\vec{\theta}_{k}-\vec{\bar{v}}\cdot k$ and $\vec{v}_{k}^{\textrm{(s)}}\coloneqq\vec{v}_{k}-\vec{\bar{v}}$, respectively. Variances are then computed for these adjusted quantities, $\vec{\theta}_{k}^{\textrm{(s)}}$ and $\vec{v}_{k}^{\textrm{(s)}}$, which show more stationary behavior. However, in Feng & Tu’s analysis, this linear movement is automatically subsumed in the first principal component due to its pronounced variance. Hence, there’s no necessity for additional subtraction of this drift in the weight covariance eigenbasis.
@@
The artifact issue becomes evident in Figure˜6, which displays synthetic data generated through stochastic gradient descent within an isotropic quadratic potential coupled with isotropic noise. With 2,500 dimensions, the model mirrors the scale of a layer in the fully connected neural network that Feng & Tu investigated. The weight series comprises 12,000 steps, which correspond to ten epochs of training this network. Analyzing this data with the weight covariance eigenbasis seemingly suggests anisotropic variance and correlation time. However, if the data is inspected without any basis change, both the variance and correlation time appear isotropically distributed as anticipated.
@@
To navigate around this key issue associated with the eigenbasis of $\vec{\Sigma}$, we adopted the eigenvectors of the Hessian matrix. Unlike $\vec{\Sigma}$, the Hessian is not computed as an average over update steps but can, in theory, be precisely calculated for any given weight vector. Consequently, the Hessian matrix does not suffer from finite size effects. The difference between these two bases for actual data is visible in Figure˜7. Here, we analyzed only the weights of the first convolutional layer of the LeNet from the main text to ensure comparability with Feng & Tu’s results. In this specific comparison, the network was trained without weight decay. Due to this and the fact that we are only investigating the weights of one layer, $\lambda_{\textrm{cross}}$ is significantly larger than all Hessian eigenvalues. As a result, when analyzing in the eigenbasis of the Hessian matrix related to this layer, both the variance and the correlation time align well with the prediction for smaller Hessian eigenvalues.
Figure 7: Comparison of weight and velocity variances for all 450 weights of the first convolutional layer of the LeNet, as discussed in the main text, analyzed in two different bases. In order to facilitate a more directly comparable analysis to f Feng & Tu [32], the network was trained without weight decay for this specific analysis and the analysis period was limited to 10 epochs, as opposed to the usual 20 epochs. The columns represent different bases: for the left column $\vec{p_{i}}$ are the eigenvectors of $\vec{\Sigma}$ and for the right column $\vec{p_{i}}$ are the eigenvectors of $\vec{H}$. The mean velocity was subtracted in the right column. The rows illustrate the weight and velocity variance, $\sigma_{\theta,i}^{2}=\vec{p_{i}}^{\top}\vec{\Sigma}\vec{p_{i}}$, $\sigma_{v,i}^{2}=\vec{p_{i}}^{\top}\vec{\Sigma_{v}}\vec{p_{i}}$ (top row), and the correlation time $\tau_{i}=2\sigma_{\theta,i}^{2}/\sigma_{v,i}^{2}$ (bottom row). The second derivative of the corresponding direction is depicted on the x-axis, $\lambda_{i}=\vec{p_{i}}^{\top}\vec{H}\vec{p_{i}}$. The top row solid lines indicate fit regions for a linear fit. For the $\vec{H}$ eigenbasis, the respective exponent of the power law relation is $1.018\pm 0.008$ for weight variance and $1.017\pm 0.002$ for velocity variance with a $2\sigma$-error. For the $\vec{\Sigma}$ eigenbasis, the corresponding exponent is $1.537\pm 0.012$ for weight variance and $1.134\pm 0.002$ for velocity variance.
However, analyzing in the eigenbasis of the weight covariance matrix, the correlation time appears heavily dependent on the second derivative of the loss in the given direction. Additionally, the relationship between the weight variance and the second derivative shifts and more closely aligns with Feng & Tu’s results, with the exponent of a power law fit being significantly larger than one. In this regime, our theory predicts proportionality between weights and Hessian eigenvalues, reflected by an exponent of one for a power law fit. The first principal component, which Feng & Tu referred to as the drift mode, stands out due to its unusually long correlation time. This is to be expected, as this is the direction in which the weights are moving at an approximately constant velocity.
G.3 Drawing with replacement
@@
Figure˜8 offers clear visual proof that when examples are selected with replacement, the previously noted anti-correlations within the SGD noise vanish. This observation confirms our hypothesis that the anti-correlations mentioned in the main text are indeed an outcome of the epoch-based learning technique. Consequently, we can predict that this change will influence the behavior of the weight and velocity variance. As previously discussed, the theoretical results we have achieved for Hessian eigenvectors with eigenvalues exceeding $\lambda_{\textrm{cross}}$ conform to what one would predict in the absence of any correlation within the noise. Therefore, when examples are drawn with replacement, we anticipate the weight variance to be isotropic in all directions, while the velocity variance should remain unchanged.
@@
Upon reviewing Figure˜9, it is clear that the velocity variance stays unchanged as predicted. However, while the weight variance remains constant for a broader subset of Hessian eigenvalues, it reduces for extremely small eigenvalues. Likewise, the correlation time is still limited for these minuscule Hessian eigenvalues. These deviations can be attributed to the finite time frame of the analysis period, comprising 20,000 update steps. This limited time window sets a cap on the maximum correlation time, consequently leading to a decreased weight variance for these small Hessian eigenvalues. Despite this, it is noteworthy that this maximum correlation time is still roughly one order of magnitude longer than the maximum correlation time induced by the correlations arising from the epoch-based learning approach.
@@
G.4 Weight variance for very small Hessian eigenvalues
As discussed in Section˜4.3, under the assumption that the gradient noise covariance is proportional to the Hessian, the stationary weight variances along Hessian eigendirections are expected to scale linearly with the corresponding eigenvalues, i.e., $\sigma_{\theta,i}^{2}\propto\lambda_{i}$ for $\lambda_{i}<\lambda_{\rm cross}$. This prediction holds well across a broad range of eigenvalues in our experiments.
However, slight deviations from this linear relationship become apparent in the lowest-curvature directions. In particular, for a subset of the smallest eigenvalues among the top 5000, the observed weight variances slightly exceed the predicted values. These deviations become more pronounced when extending the analysis to the top 10,000 eigenvalues, as shown in Figure˜10.
Figure 10: Relationship between Hessian eigenvalues and the variances of weights and velocities. The solid lines signify theoretical predictions from Equation˜6, assuming $\boldsymbol{\mathbf{C}}\approx c_{0}\boldsymbol{\mathbf{H}}$ with $c_{0}$ fitted to the data. The plot extends the analysis to the top 10,000 Hessian eigenvectors. While the predicted proportionality holds well over a broad range, slight deviations appear in the lowest-curvature directions, where some weight variances begin to saturate and no longer decrease with $\lambda_{i}$. These deviations are consistent with either non-equilibrated weights in flat directions or corrections to the noise anti-correlations due to fluctuating weights.
This behavior may arise from several factors. First, modifications to the noise anti-correlations due to a fluctuating weight vector can influence the linear variance–curvature relationship, particularly for very small Hessian eigenvalues. In ˜E, we discussed an extended model that accounts for such fluctuations, predicting a minimal floor for the weight variance as $\lambda_{i}\to 0$.
Another possibility is that the weights have not yet fully equilibrated in these extremely flat directions within the training duration considered, leading to residual variances above the predicted values. The smallest eigenvalues observed are on the order of $10^{-1}$. To illustrate the plausibility of slow equilibration, consider full gradient descent with the final learning rate of about $10^{-5}$. Even with momentum $\beta=0.9$, it would take roughly $10^{5}$ update steps to make substantial progress in such directions. In our setup, this corresponds to around 100 epochs – while training lasted only 300 epochs in total. Using SGD instead of full-batch GD likely increases this timescale further due to additional noise.
Despite these deviations, the overall trend of variance–curvature proportionality remains robust across a wide range of eigenvalues. Where deviations do occur, the weight variance is still significantly lower than predictions from prior literature, which assumed a uniform variance across all directions equal to that in the sharpest mode.
G.5 Testing different Hyperparameters
In this section we examine the dependence of the theoretical predictions on the three hyperparameters learning rate $\eta$, momentum $\beta$ and batch size $S$. For this, we train the LeNet again for 300 epochs, using an exponential learning rate schedule that reduces the learning rate by a factor of $0.98$ every epoch and afterwards we perform the numerical analysis as described in the main text.
However, we now train the network several times, always varying one of the hyperparameters while keeping the other two fixed. If not varied, the momentum was set to 0.90 and the batch size was set to 64. To ensure that training always achieves a comparatively high training accuracy, the initial learning rate was set to 0.005 when the batch size is varied and to 0.02 when the momentum is varied. Five different values are examined for each hyperparameter. To investigate the dependencies on the learning rate, the values 0.005, 0.01, 0.02, 0.03, and 0.04 were used for training. For momentum, the values 0.00, 0.50, 0.75, 0.90, and 0.95 were examined, and for batch size, the values 32, 50, 64, 100, and 128 were examined.
@@ G.5 Testing different Hyperparameters
Figure˜11 shows as an example the weight and velocity variances as well as the correlation times for different values of the batch size for one training seed each. It can be seen that the theory is not only valid for the hyperparameter combination from the previous section, but is also generally applicable for different hyperparameters. In particular, we see that there is still good agreement with the theory even if the strictly necessary condition for the theoretical derivation of the noise autocorrelation, that the number of batches per epoch $M=N/S$ is an integer, is not met.
@@
For the evaluation of the dependence of these variables on the hyperparameters, they were determined as follows for the various hyperparameter combinations using the data from the respective correlation time plot. For the maximum correlation time $\tau_{\textrm{SGD}}$, the average of all correlation times was taken for which the corresponding Hessian eigenvalue is smaller than the theoretical crossover value. However, the result for the numerically determined maximum correlation time is similar when taking the average of all determined correlation times for a hyperparameter combination, since only very few Hessian eigenvalues are larger than the crossover value.
@@
For the numerical determination of the crossover value $\lambda_{\rm cross}$, a linear function was first fitted to the correlation times of the 5 largest Hessian eigenvalues in the log-log plot of the correlation times against the Hessian eigenvalues. In this region of the first 5 values, the correlation time is always clearly dependent on the Hessian eigenvalue and does not yet belong to the region of constant correlation times. The intersection of the fitted line with the numerically determined maximum correlation time $\tau_{\textrm{SGD}}$ is then taken as the crossover value $\lambda_{\rm cross}$. If the numerically determined correlation times follow the theory exactly, then the correlation times determined in this way for $\tau_{\textrm{SGD}}$ and $\lambda_{\rm cross}$ would also follow the theory accurately.
And indeed, Figure˜12 shows a good agreement between the theory and the numerically determined values, although it should be noted that the deviations are larger than the random fluctuations between the different seeds.
G.6 Autocorrelation of SGD noise at the beginning of training
The theoretical prediction for the autocorrelation of SGD noise given by Equation˜2 was derived for a static weight vector and Figure˜1 shows strong agreement with empirical data from later stages of training. However, further numerical investigations reveal, that even at the beginning of training, when the network weights are clearly changing significantly, SGD noise still exhibits anti-correlations. Figure˜13 demonstrates that these pronounced anti-correlations are also in good agreement with our theoretical prediction, although our analytic theory relates only to a late phase of training.
Figure 13: Autocorrelations of the SGD noise for the first ten epochs of the LeNet training schedule described in the main text. Since the Hessian matrix has a higher variability at this stage of training compared to later stages and its eigenvectors are not as suitable, for this experiment only, the noise terms are projected onto the first 2,500 principal components of the weight trajectory instead of the Hessian eigenvectors. As no variances are analyzed, no significant artifacts due to finite size effects are expected. The autocorrelation of the projected noise terms is computed individually and then averaged over the 2,500 principal components.
G.7 Relation between Hessian and noise covariance